# Unveiling Super Experts in Mixture-of-Experts Large Language Models

### v1

@@ Abstract
Sparsely activated Mixture-of-Experts (MoE) models have shown promise in enhancing the learning capacity of large language models (LLMs), yet their substantial number of parameters presents significant challenges for the deployment of MoE LLMs. Leveraging the intrinsic importance differences among experts, recent research has explored expert-level compression techniques to improve the efficiency of MoE LLMs. However, existing approaches often rely on empirical criteria to identify critical experts, lacking a deeper exploration and understanding of the heterogeneous importance of experts. In this study, we present the first discovery and investigation of a distinct subset of experts that play a crucial role in the underlying mechanisms during the model’s forward inference. These experts are prevalent in open-source MoE LLMs, and despite their limited number, pruning them leads to a significant decline in model performance (e.g., pruning three causes Qwen3-30B-A3B to produce repetitive and uninformative outputs). We refer to these experts as Super Experts (SEs). Our comprehensive analysis provides progressively deeper insights into SEs. (i) SEs are characterized by rare but extreme activation outliers in the output of the down_proj, which give rise to massive activations in the hidden states between decoder layers. Moreover, the distribution of SEs remains model-specific and is unaffected by post-training processes. (ii) By pruning SEs, we assess their significance across a variety of tasks, revealing their considerable impact on the model’s overall performance, particularly in mathematical reasoning. (iii) We further enhance our understanding of the influence of SEs compression. Our findings confirm that MoE LLMs rely on SEs to induce attention sinks, which are crucial for the distribution of attention scores but are significantly disrupted by SE pruning. We also develop an automated tool for the rapid and precise profiling of SEs in new MoE LLMs. The code is available at https://github.com/ZunhaiSu/Super-Experts-Profilling.
@@
1 Introduction
Sparsely activated Mixture-of-Experts (MoE) models employ dynamic routing and sparse activation, demonstrating significant potential in enhancing the learning capacity of large language models (LLMs) (Cai et al. 2024; Mu and Lin 2025). This paradigm has led to the development of state-of-the-art MoE LLMs, including DeepSeek (Guo et al. 2025; Liu et al. 2024b), Qwen (Team 2024b; Yang et al. 2025), and others (Meta 2025; Jiang et al. 2024). Despite their potential, a significant challenge with MoE LLMs stems from their large parameter size and high computational cost (Li et al. 2023; Lu et al. 2024; Chowdhury et al. 2024), which present considerable obstacles for deployment. Model compression techniques, such as quantization (Frantar et al. 2022; Xiao et al. 2023a; Su et al. 2025) and pruning (Frantar and Alistarh 2023; Sun et al. 2023; Ma, Fang, and Wang 2023), enable the development of more compact and computationally efficient models (Zhu et al. 2024; Wang et al. 2024).
Beyond LLM-oriented compression approaches, expert-level compression methods have been developed by leveraging the structural characteristics of MoE models and the uneven importance of experts induced by training strategies (Chowdhury et al. 2024; Chi et al. 2022; Lu et al. 2024). Specifically, expert-level compression employs various expert importance metrics to guide the pruning, merging, or skipping of less critical experts (Lu et al. 2024; Huang et al. 2025; Xie et al. 2024), prioritize more important ones by assigning higher bit budgets during quantization (Duanmu et al. 2025; Li et al. 2024; Zheng et al. 2025), and allocate more ranks in low-rank decomposition (Yang et al. 2024; Li et al. 2023). For instance, several studies evaluate expert importance based on access frequency or other statistical features derived from the routers in MoE layers (Chen et al. 2022; Li et al. 2024, 2023; Huang et al. 2025).
@@ 1 Introduction
Analyzing expert importance not only facilitates model compression but also provides deeper insights into the inner workings of MoE LLMs. However, existing approaches often rely on empirical criteria to identify critical experts, lacking a deeper exploration and understanding of the heterogeneous importance among experts. In this study, we address a fundamental yet previously overlooked question: Is there a distinct subset of experts that plays a critical role in the underlying mechanisms during the forward inference of MoE LLMs? Through comprehensive analysis of various open-source MoE LLMs, we consistently confirm the existence of such experts. Despite their limited number, pruning these experts leads to a significant collapse in model performance. As shown in Figure 1, pruning just three experts from Qwen3-30B-A3B (Yang et al. 2025) leads to a significant degradation in model performance, while randomly pruning other experts results in a considerably smaller impact. We refer to these experts as Super Experts (SEs), and our subsequent comprehensive analysis offers increasingly deeper insights into SEs.
@@
In Section 3, we first characterize SEs and explore their distribution across different models and input data domains. SEs are identified by rare but extreme activation outliers in the output of the down_proj, which induce massive activations (Sun et al. 2024) in the hidden states between decoder layers. Intriguingly, the distribution of SEs remains model-specific, and the SEs in the base model maintain consistency after post-training processes. For instance, Qwen3-30B-A3B-Base and Qwen3-30B-A3B exhibit identical SEs distribution patterns. In Section 4, we assess the importance of SEs by quantifying performance degradation across multiple capabilities following their pruning. For non-reasoning models, all few-shot evaluations we conducted demonstrate considerable performance degradation, with the most pronounced decline observed in the GSM8K dataset (Cobbe et al. 2021). Notably, for reasoning models, pruning SEs leads to a complete performance collapse, with Pass@1 dropping to nearly zero on tasks such as AIME and Math-500 (AIME 2024, 2025; Lightman et al. 2023). This highlights the critical role of SEs in the model’s overall performance, particularly in mathematical reasoning tasks.
In Section 5, we further enhance our understanding of the influence of SE compression. Our findings confirm that MoE LLMs rely on SEs to induce attention sinks, which are crucial for the distribution of attention scores and must be preserved during sparse attention or token compression (Xiao et al. 2023b; Su et al. 2025). We quantitatively assess the significant disruption of attention sinks caused by SE compression, as measured by the proposed attention sink decay rate. Our contributions are summarized as follows:
  • We present the first discovery and investigation of Super Experts (SEs) in MoE LLMs. We identify and localize SEs across various prominent open-source MoE LLMs, thoroughly validating that their distribution remains model-specific and unaffected by post-training processes. Additionally, we develop an automated tool for rapid and precise profiling of SEs in new models.
  • We conduct extensive experiments to evaluate the importance of SEs, revealing their significant impact on the model’s overall performance, particularly in mathematical reasoning abilities.
  • We further enhance the understanding of SEs compression. Our findings confirm that MoE LLMs rely on SEs to induce attention sinks. Using the proposed attention sink decay rate, we quantitatively assess the significant disruption of attention sinks caused by SE pruning.
  • This discovery of SEs provides new insights into the inner workings of MoE LLMs and highlights the critical importance of preserving SEs during model compression. These insights offer valuable guidance for achieving more effective MoE LLM compression strategies.
Figure 2: Decoder Architecture of MoE LLM.
@@ 2 Preliminaries on MoE LLMs
MoE LLMs
LLMs are typically structured as a stack of Transformer decoder blocks (Vaswani et al. 2017), each consisting of a multi-head self-attention (MHSA) layer and a feed-forward network (FFN) layer. In MoE LLMs, the FFN layers are replaced by MoE layers, where each layer consists of multiple experts, each represented by a FFN. A concise overview of MoE LLMs is presented in Figure 2. Let $H^{0}\in\mathbb{R}^{n\times d}$ represent the input to the first decoder, where $d$ is the embedding dimension, and $n$ is the length of the tokenized input sequence. Then, the output of the $l$-th decoder block, $H^{l}\in\mathbb{R}^{n\times d}$, is given by:
@@
Figure 3: Super Experts mechanism in Qwen3-30B-A3B. The line plots depict the maximum output magnitudes of down_proj for experts 68/92/82 across layers. Massive activation is gradually amplified through expert 68 in layer 1, expert 92 in layer 2, and expert 82 in layer 3. Extreme activation outliers from these Super Experts are propagated into the hidden states between decoders via residual summation, progressively leading to massive activation.
Figure 4: Impact of SEs pruning on massive activations in Qwen3-30B-A3B. Massive activations are computed using 100 input samples from the C4 (Raffel et al. 2020) dataset, each with a length of 2K.
Qwen3-30B-3B 6144 3 0.05% 744.0 540.0 430.0 63.5 19.1 12.1 7.3 74.4
DeepSeek-V2-Lite-Chat 1782 2 0.22% 1424.0 462.0 112.5 89.5 37.5 24.0 34.5 142.4
Table 1: The activations identified as Super Experts are highlighted in bold.
MHSA Layer
After LN, the input $H^{l-1}$ is projected through the weight matrices $W^{l}_{Q},W^{l}_{K},W^{l}_{V}\in\mathbb{R}^{d\times d}$ to generate the Query, Key, and Value, which are then divided into $K$ heads, denoted as $Q^{l,k},K^{l,k},V^{l,k}$, for $1\leq k\leq K$. The MHSA is computed as:
$A^{l,k}=\text{Softmax}\left(\frac{Q^{l,k}K^{l,k^{T}}}{\sqrt{d_{k}}}+M\right),$ (3)
$\quad O^{l}=\text{Concat}_{k=1}^{K}\left(A^{l,k}V^{l,k}\right)W^{l}_{O},$ (4)
where $M$ represents the attention mask, and $d_{k}=d/K$. For simplicity, the rotation position encoding (RoPE) (Su et al. 2024) is omitted here.
MoE Layer
Next, $H^{l^{\prime}}$ is passed through a new LN and enters the MoE layer. First, the router network determines which experts to activate and how to scale their outputs through the weight matrix $W_{G}$. The routing weights $G\in\mathbb{R}^{n\times E}$ are computed as:
@@
[…]
@@
[…]
@@
(b) DeepSeek-V2-Lite-Chat.
(c) Mixtral-8x7B-v0.1-instruct.
(d) Qwen3-30B-A3B-Base.
(e) DeepSeek-V2-Lite.
@@
Figure 5: Heatmap visualizations of the maximum output magnitudes from the down_proj for each expert across layers. The analysis is conducted using the C4 (Raffel et al. 2020) dataset. SEs are highlighted with arrows.
3 Super Experts: Discovery and Localization
In this section, we first demonstrate the discovery process of SEs using Qwen’s latest MoE LLM, Qwen3-30B-A3B, as an example. Next, we analyze SEs across different MoE LLMs and data domains to examine their distribution patterns and highlight the widespread presence of SEs.
3.1 Exploring Super Experts
Massive Activations in MoE LLMs
The discovery of SEs arises from an exploration and analysis of the formation of massive activations (Sun et al. 2024) in MoE LLMs. Recent research has explored a distinct class of extreme activation outliers in LLMs, which appear in the hidden states between decoder layers and are known as massive activations (Sun et al. 2024; Guo et al. 2024; Yu et al. 2024). They are limited in number, yet their values are orders of magnitude larger than those of other activations (e.g., up to 100,000 times larger). Existing research has yet to extensively explore massive activations in MoE LLMs and the mechanisms behind their formation. To address this gap, we first validate the existence of massive activations across several open-source MoE LLMs. Our findings reveal that massive activations persist in all MoE LLMs we investigated, as shown in the first image on the left in Figure 4.
Super Experts Induce Massive Activations
Given that massive activations are also present in MoE LLMs, a deeper question arises: How are these massive activations formulated in MoE LLMs? Are they the result of all activated experts working together, or are they driven by only a few specific experts, or perhaps caused by other parts of the model? Through observations of several prominent open-source MoE LLMs (e.g., Qwen series, DeepSeek series, Mistral), we surprisingly find that a small subset of experts consistently generates rare yet extreme activation outliers in the output of their down_proj layers. These outliers are subsequently passed onto the hidden states via residual summation after the MoE layers, leading to massive activations. The entire process is illustrated in Figure 3 using the Qwen3-30B-A3B example. This phenomenon typically occurs in a single layer (e.g., Mixtral) or in just a few layers (e.g., Qwen3-30B-A3B) starting from the initial decoder layers, ultimately leading to stable massive activations across nearly all subsequent layers. To directly validate this mechanism, we also perform ablation experiments by pruning the SEs in Qwen3-30B-A3B. As illustrated in Figure 4, pruning SEs from a single layer effectively eliminates their contribution to massive activations. Furthermore, when all SEs are pruned, massive activations are completely eliminated, confirming that they are directly generated by SEs.
3.2 Localization of Super Experts
In this section, we begin by discussing how to accurately identify SEs and then examine their distribution patterns across various models and data domains.
Super Experts Profiling
Given that SEs are defined by their influence on the formation of massive activations through the extreme activation outliers they generate, we propose the following broad yet effective quantitative definition. Specifically, we compute the maximum output magnitudes to the down_proj for all experts across all layers. Let $L$ denote the set of layers responsible for the formation of massive activations. Let $a_{l,e}$ denote the maximum output magnitude to the down_proj of expert $e$ in layer $l$, and let $\mathcal{A}=\{a_{l,e}\}$ be the set of all such values across the entire model. An expert $e$ in layer $l$ is classified as a SE if:
where $P_{99.5}=\text{Percentile}_{99.5}(\mathcal{A})$ and $a_{\text{max}}=\max\mathcal{A}$. This criterion effectively identifies the experts of interest across various MoE LLMs we investigated, as highlighted in bold in Table 1. We have developed an automated tool for rapid and precise SE profiling based on this definition. The code is available at https://github.com/ZunhaiSu/Super-Experts-Profilling.

### v2

@@ Abstract
Leveraging the intrinsic importance differences among experts, recent research has explored expert-level compression techniques to enhance the efficiency of Mixture-of-Experts (MoE) large language models (LLMs). However, existing approaches often rely on empirical heuristics to identify critical experts, while lacking a deeper understanding into the heterogeneous importance of experts and the inner workings of MoE LLMs. In this study, we report, for the first time, the discovery and systematic investigation of a distinct subset of experts that play a pivotal role in the model’s forward inference. These experts are prevalent in open-source MoE LLMs, and despite their extremely limited number, pruning them results in a substantial decline in model performance (e.g., prune just three out of 6,144 causes Qwen3-30B-A3B to generate repetitive and uninformative outputs). We refer to these experts as Super Experts (SEs). Our comprehensive analysis provides progressively deeper insights into SEs: (i) SEs are characterized by rare but extreme activation outliers in the output of the down_proj, which give rise to massive activations in the hidden states between decoder layers. Moreover, the distribution of SEs is model-specific, data-agnostic, and remains unaffected by post-training processes. (ii) By pruning SEs, we assess their significance across a variety of tasks, revealing their considerable impact on the model’s overall performance, particularly in mathematical reasoning. (iii) We further investigate why compressing SEs exerts such a pronounced impact. We show that, in MoE LLMs, SEs serve as the primary source of the systematic outlier mechanism in Transformers, and that compressing them profoundly disrupts this process, ultimately causing the collapse of attention sinks. These findings advance the understanding of the internal dynamics of MoE LLMs, filling an important gap in the current knowledge. In addition, we developed an automated tool for rapid and accurate SE profiling. The code is provided in https://github.com/ZunhaiSu/Super-Experts-Profilling.
1 Introduction
Sparsely activated Mixture-of-Experts (MoE) models employ dynamic routing and sparse activation, demonstrating significant potential in enhancing the learning capacity of large language models (LLMs) (Cai et al., 2024; Mu & Lin, 2025). This paradigm has led to the development of state-of-the-art MoE LLMs, including DeepSeek (Guo et al., 2025; Liu et al., 2024b), Qwen (Yang et al., 2025a), LongCat-Flash (Team et al., 2025) and others. Despite their potential, a significant challenge stems from their large parameter size and high computational cost (Li et al., 2023; Lu et al., 2024; Chowdhury et al., 2024), which present considerable obstacles for deployment. Model compression techniques, such as quantization (Frantar et al., 2022; Xiao et al., 2023a; Su et al., 2025a), pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and others (Zhu et al., 2024; Wang et al., 2024), enable the development of more compact and computationally efficient models.
Beyond LLM-oriented compression approaches, expert-level compression methods have been developed by leveraging the structural characteristics of MoE models and the uneven importance of experts induced by training strategies (Chowdhury et al., 2024; Chi et al., 2022; Lu et al., 2024). Specifically, it employs various expert importance metrics to guide the pruning, merging, or skipping of less critical experts (Lu et al., 2024; Huang et al., 2025; Xie et al., 2024), prioritize more important ones by assigning higher bit budgets during quantization (Duanmu et al., 2025; Li et al., 2024), and allocate more ranks in low-rank decomposition (Yang et al., 2024a; Li et al., 2023).
@@ 1 Introduction
For instance, several works evaluate expert importance by measuring activation frequency or by analyzing router scores produced within MoE layers (Li et al., 2024; 2023; Huang et al., 2025). Additionally, reconstruction loss and other similarity-based metrics have been utilized in previous studies (Lu et al., 2024; Duanmu et al., 2025; Zhang et al., 2024b).
@@
Analyzing expert importance not only facilitates model compression but also provides deeper insights into the inner workings of MoE LLMs. However, existing approaches often rely on empirical criteria to identify critical experts, lacking a deeper exploration and understanding of the heterogeneous importance among experts. In this study, we address a fundamental yet previously overlooked question: Is there a small subset of distinct experts that plays an exceptionally critical role in the underlying mechanisms of MoE LLMs?
Through comprehensive analysis of various open-source MoE LLMs, we consistently confirm the existence of such experts. Despite their extreme limited number, pruning these experts leads to a significant collapse in model performance. As shown in Figure 1, pruning just three experts from Qwen3-30B-A3B leads to a significant degradation in model performance, while randomly pruning other experts results in a considerably smaller impact. We refer to these experts as Super Experts (SEs), and our comprehensive analysis provides progressively deeper insights into SEs.
@@
In Section 3, we first characterize SEs and analyze their distribution across various models and input data domains. SEs are identified by extreme activation outliers in the output of the down_proj, which induce massive activations (MAs) (Sun et al., 2024). Intriguingly, the distribution of SEs remains model-specific, data-agnostic, and the SEs in the base model maintain consistency after post-training processes. In Section 4, we assess the importance of SEs by quantifying performance degradation following their dynamic pruning. Notably, pruning SEs leads to a complete performance collapse with Pass@1 dropping to nearly zero on tasks such as AIME and Math-500 (AIME, 2024; 2025; Lightman et al., 2023) for reasoning LLMs. In Section 5, we further deepen our understanding of SEs by revealing their origin in the behavior of systematic outliers mechanism in Transformers (Su & Yuan, 2025; An et al., 2025). Our findings confirm that MoE LLMs rely on SEs to induce attention sinks, which are crucial for the distribution of attention scores and must be preserved during sparse attention or KV compression (Xiao et al., 2023b; Su et al., 2025a).
The main contributions of this work are summarized as follows:
$\bullet$ We provide the first comprehensive characterization of Super Experts (SEs) in MoE LLMs, an exceptionally rare yet fundamentally critical subset of experts, thereby filling a significant gap in the existing understanding of MoE LLMs. Extensive analyses across multiple models and tasks reveal key properties of SEs, including their stable distribution and critical impact on model performance.
$\bullet$ We demonstrate that SEs are the primary drivers of systematic outliers in Transformers. In MoE LLMs, their strong activation on attention sink tokens makes them the fundamental source of these outliers, and compressing them severely disrupts this process, ultimately leading to the collapse of attention sinks.
$\bullet$ Our findings on SEs provide new insights into the internal dynamics of MoE LLMs and the heterogeneous importance of experts. These insights serve as a foundation for designing more expert-balanced pre-training regimes and for advancing robust expert compression strategies.
@@ 2 Preliminaries on MoE LLMs
MoE LLMs. LLMs are typically structured as a stack of Transformer decoder blocks (Vaswani et al., 2017), each consisting of a multi-head self-attention (MHSA) layer and a feed-forward network (FFN) layer. In MoE LLMs, the FFN layers are replaced by MoE layers, where each layer consists of multiple experts, each represented by a FFN. A concise overview of MoE LLMs is presented in Figure 2. Let $H^{0}\in\mathbb{R}^{n\times d}$ represent the input to the first decoder, where $d$ is the embedding dimension, and $n$ is the length of the tokenized input sequence. Then, the output of the $l$-th decoder block, $H^{l}\in\mathbb{R}^{n\times d}$, is given by:
@@
Figure 2: Decoder Architecture of MoE LLM.
MoE Layer. The hidden representation after MHSA, $H^{l^{\prime}}$, passes through a LN and then fed into the MoE layer. First, the router network determines which experts to activate and how to scale their outputs through the weight matrix $W_{G}$. The routing weights $G\in\mathbb{R}^{n\times E}$ are computed as:
@@
Then, sparse activation of the experts is achieved by selecting the top-$k$ routing weights for each input token. The output of the activated experts is scaled by the routing weights and aggregated to form the output of the MoE layer:
@@
where $\text{Top-$k$}(G_{j})$ denotes the indices of the top-$k$ routing weights for the $j$-th input token. The FFN is defined as:
@@
3 Super Experts: Discovery and Localization
In this section, we first demonstrate the discovery process of SEs using Qwen3-30B-A3B as an example. Next, we analyze SEs across different MoE LLMs and data domains to examine their distribution patterns and highlight the widespread presence of SEs.
3.1 Super Experts Induce Massive Activations
Recent research has explored a distinct class of extreme activation outliers in LLMs, which appear in the hidden states between decoder layers and are known as massive activations (MAs) (Sun et al., 2024; Guo et al., 2024). They are limited in number, yet their values are orders of magnitude larger than those of other activations (e.g., up to 100,000 times larger). The discovery of SEs arises from an exploration and analysis of the formation of MAs (Sun et al., 2024) in MoE LLMs. Existing research has yet to clarify how these MAs arise in MoE LLMs. Do these activations arise from the collective activity of all activated experts, are they primarily driven by some specific experts, or are they instead caused by other components of the model?
Through analysis of several prominent open-source MoE LLMs (e.g., Qwen series, DeepSeek series, Mistral), we surprisingly find that a small subset of experts consistently produces extreme activation outliers in the output of their down_proj layers. These outliers are subsequently passed onto the hidden states via residual summation after the MoE layers, leading to MAs. The entire process is illustrated in Figure 3 using Qwen3-30B-A3B as example. This phenomenon typically occurs in a single layer (e.g., Mixtral) or in just a few layers (e.g., Qwen3-30B-A3B) starting from the initial decoder layers, ultimately leading to stable MAs across nearly all subsequent layers. To directly validate this mechanism, we also perform ablation experiments by dynamic pruning the SEs in Qwen3-30B-A3B. As illustrated in Figure 4, pruning SEs from a single layer effectively eliminates their contribution to MAs. Furthermore, when all SEs are pruned, MAs are completely eliminated, confirming that they are directly generated by SEs.
3.2 Localization of Super Experts
Figure 3: SEs mechanism in Qwen3-30B-A3B. The line plots depict the maximum output magnitudes of down_proj for experts 68/92/82 across layers. Massive activation is gradually amplified through expert 68 in layer 1, expert 92 in layer 2, and expert 82 in layer 3. Extreme activation outliers from these SEs are propagated into the hidden states between decoders via residual summation, progressively leading to massive activation.
Figure 4: Impact of SEs pruning on MAs in Qwen3-30B-A3B. MAs are computed using 100 input samples from the C4 (Raffel et al., 2020) dataset, each with a length of 2K.
Table 1: Activations identified as SEs are highlighted in bold, based on results from the C4 dataset.
Qwen3-30B-A3B 6144 3 0.05% 744.0 540.0 430.0 63.5 19.1 12.1 7.3 74.4
DeepSeek-V2-Lite-Chat 1782 2 0.11% 1424.0 462.0 112.5 89.5 37.5 24.0 34.5 142.4
3.2.1 Super Experts Profiling
Given that SEs are defined by their influence on the formation of MAs through the extreme activation outliers they generate, we propose the following broad yet effective quantitative definition. Specifically, we compute the maximum output magnitudes to the down_proj for all experts across all layers. Let $L$ denote the set of layers responsible for the formation of MAs. Let $a_{l,e}$ denote the maximum output magnitude to the down_proj of expert $e$ in layer $l$, and let $\mathcal{A}=\{a_{l,e}\}$ be the set of all such values across the entire model. An expert $e$ in layer $l$ is classified as a SE if:
where $P_{99.5}=\text{Percentile}_{99.5}(\mathcal{A})$ and $a_{\text{max}}=\max\mathcal{A}$. This criterion effectively identifies the experts of interest across various MoE LLMs we investigated, as highlighted in bold in Table 1. No specific dataset is designated for identifying SEs, since we later demonstrate that their distribution remains stable across different input datasets. The pseudocode of SEs profiling is presented in Appendix I. We have developed an automated tool for rapid and precise SE profiling based on this definition. The code is provided in https://github.com/ZunhaiSu/Super-Experts-Profilling.
@@ 3.2.1 Super Experts Profiling
(b) Qwen3-30B-A3B-Base.
(c) DeepSeek-V2-Lite-Chat.
(d) DeepSeek-V2-Lite.
(e) Mixtral-8x7B-v0.1-instruct.
@@
Figure 5: Heatmap visualizations of the maximum output magnitudes from the down_proj for each expert across layers. SEs are highlighted with arrows.
3.2.2 Distribution of Super Experts Across Models and Data Domains
We select three representative MoE LLMs with distinct designs for analysis: Qwen3-30B-A3B, DeepSeek-V2-Lite-Chat, and Mixtral-8x7B-Instruct-v0.1. We also include the base models of these three LLMs to illustrate the impact of post-training processes. Although all of these models are MoE LLMs, they exhibit distinct design differences.
Table 2: SEs of several MoE LLMs.